# Efficient Conditional Gradient Methods for Solving Stochastic Convex Bilevel Optimization Problems

### v1

@@ 3.2 Minimal norm logistic regression
IR-SCG $1726604.0$ $1726604.0$
IR-FSCG $225263.8$ $19926327.2$
SBCGI $6207.8$ $8852217.8$
SBCGF $5986.5$ $9376410.0$
aR-IP-SeG (short step) $1218831.0$ $2437662.0$
aR-IP-SeG (long step) $973268.7$ $1946537.4$
SDBGD (short step) $1623034.9$ $1623034.9$
SDBGD (long step) $1618114.5$ $1618114.5$
Figure 2: The inner-level optimality gap (left) and the outer-level absolute optimality gap (right) over time for different algorithms on the minimal norm logistic regression problem Eq. 8.
Appendix A Convergence proofs
This section provides the proofs of Theorems 1 and 2. We employ the notation $\Phi_{t}(x):=\sigma_{t}F(x)+G(x)$ and $\widehat{\nabla\Phi}_{t}:=\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}$ to denote the gradient estimate at step $t$. All norms refer to the Euclidean norm.
A.1 Preparatory lemmas
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with some given stepsizes $\{\alpha_{t}\}_{t\geq 0}$. If 1 holds, then for every $t\geq k$, we have
$\Phi_{t}(x_{t+1})-\Phi_{t}(x_{\operatorname{opt}})\!\leq\!(1-\alpha_{t})(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}}))+\frac{(L_{f}\sigma_{t}\!+\!L_{g})D^{2}\alpha_{t}^{2}}{2}+2\alpha_{t}D\left\|\nabla\Phi_{t}(x_{t})\!-\!\widehat{\nabla\Phi}_{t}\right\|.$
Proof of Lemma 1.
@@
$x_{t+1}=x_{t}+\alpha_{t}(v_{t}-x_{t}),\quad\text{where}\quad v_{t}\in\operatorname*{arg\,min}_{v\in X}v^{\top}\widehat{\nabla\Phi}_{t}.$
@@
Let $x_{\operatorname{opt}}\in X_{\operatorname{opt}}$, and set $w_{t}=x_{\operatorname{opt}}$. Thanks to the convexity of $\Phi_{t}$, one can easily show that $\nabla\Phi_{t}(x_{t})^{\top}(x_{\operatorname{opt}}-x_{t})\leq\Phi_{t}(x_{\operatorname{opt}})-\Phi_{t}(x_{t})$. Thus, we obtain
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$. If 1 holds, then for each $t\geq k$, we have
@@
Applying Lemma 1 with stepsize $\alpha_{t}=2/(t+2)$ for all $t\geq k$ gives the recursion $h_{t+1}\leq\frac{t}{t+2}h_{t}+\eta_{t}$ for any $t\geq k$. This can be re-expressed as
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 or Algorithm 2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$. If 1 holds, then for each $t\geq k$, we have
[…]
@@
Proof of Proposition 1.
By definition of the function $\Phi_{t}$, the first term in Eq. 10 satisfies
@@
By construction, $x_{0},v_{t}\in X$ Hence, $x_{t+1}\in X$ as it is a convex combination of $x_{t}$ and $v_{t}$, and $X$ is assumed to be convex. By a simple induction argument, one can easily show that the sequence $\{x_{t}\}_{t\geq 0}\in X$. Therefore, we have $\min_{x\in X}F(x)-F_{\operatorname{opt}}\leq F(x_{t})-F_{\operatorname{opt}}$ for all $t\geq 0$. Using this bound together with the identity in Eq. 13, we can further lower bound the left-hand side of inequality (10), and arrive at the bound (11). This completes the proof of the first claim.
For the second claim, since $\{x_{t}\}_{t\geq 0}\in X$ is feasible in the lower-level problem, it follows directly that $G(x_{t})\geq G_{\operatorname{opt}}$. Combining this inequality with (13) and  (10), we obtain Eq. 12, which concludes the proof. ∎
A.2 Proof of Theorem 1
The proof proceeds by applying Proposition 1 with $k=0$. In this case, the main random component in Eq. 11 and Eq. 12 is the term $\sum_{i\in[t]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$. We first derive a probabilistic upper bound for this term, then carefully set the parameters in Algorithm 1 to establish the convergence rate for both the lower- and upper-level problems.
We begin by bounding the gradient estimator used in Algorithm 1.
@@ A.2 Proof of Theorem 1
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with stepsize $\alpha_{t}=2/(t+2)$ for any $t\geq 0$. If 1 holds, then for any $t\geq 1$, given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$, it (jointly) holds that
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$. If 1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$ and any $t\geq 1$, we have
$\sum_{i\in[t]}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|\leq c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)(t+1)^{3/2}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}.$
Proof of Proposition 2.
Given $i\geq 1$ and $\delta\in(0,1)$, by Lemma 3, with probability at least $1-\delta/i(i+1)$, we have
@@
To establish the desired convergence results for Algorithm 1, we need to impose certain conditions on the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$, stated below.
@@
The regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing, positive, and converge to 00.
@@
While it is not obvious a priori, Lemma 16 in Section B.4 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 3. In particular, 1 to 3 hold when $\sigma_{t}=\varsigma(t+1)^{-p}$ for any $p\in(0,1)$, which is what is used in Theorems 1 and 2, and Lemma 6 provides explicit bounds for this case. We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm 1 for the inner-level problem.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta}$ be defined as in (15a). If 1 and 1 to 3 hold, with probability at least $1-\delta$, it holds that
@@
Proof of Lemma 4.
Combining (11) from Proposition 1 with the probabilistic bound in Proposition 2, and after some straightforward calculations, with probability at least $1-\delta$, we have
@@
We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm 1 for both outer- and inner-level problems.
@@
Let $\{z_{t}\}_{t\geq 1}$ denote the iterates generated by Algorithm 1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta},V$ be defined as in Eq. 15. If 1 and 1 to 3 hold, then with probability at least $1-\delta$, it (jointly) holds that
@@
Proof of Lemma 5.
Recall from Algorithm 1 that we have defined
@@
Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to 1, and applying Proposition 2, with probability at least $1-\delta$, we have
@@
Lemma 5 establishes a convergence result in terms of the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$. The next lemma specifies an update rule for these parameters and provides bounds on the quantities introduced in Eq. 15.
@@
Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq. 15.
@@
    If $p\in(0,1)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies 1 and 2 with $L=p$. Furthermore,
@@
    If $p\in(0,1/2)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies 3, and we have
    $\bar{C}_{\delta}\leq 4c\left(8\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left(\frac{8d}{\delta}\right)}.$
Proof of Lemma 6.
As for assertion (i), it is trivial to see that the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies Conditions 1. To validate 2, observe that
@@
Dividing both sides by $t(t+1)\sigma_{t}^{2}$, we deduce that $V\leq 2p$. When $01-2p<$ 0, we have
@@
As for assertion (ii), it is trivial to see that 3 holds. We thus focus on bounding $\bar{C}_{\delta}$. Using the fact that $\sup_{t\geq 1}{\log(t)}/{t}={1}/{e},$ for any $t\geq 1$, we may conclude that
@@
$=8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{1}{(t+1)^{1-2p}}\left(\frac{2}{1-2p}\log\left(t^{1-2p}\right)+\log\left(\frac{8d}{\delta}\right)\right)}$
$\leq 8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left(\frac{8d}{\delta}\right)}.$

### v2

@@ 4.2 Dictionary learning
IR-SCG $32736.1$ $65472.2$
IR-FSCG $30532.6$ $1872851.0$
SBCGI $466.3$ $7440.6$
SBCGF $361.84$ $6147916.0$
aR-IP-SeG (short step) $6290.0$ $25160.0$
aR-IP-SeG (long step) $6278.1$ $25112.4$
SDBGD (short step) $11441.3$ $22882.6$
SDBGD (long step) $12239.3$ $24478.6$
Appendix A Preparatory Lemmas
Throughout all appendices, we employ the notation $\Phi_{t}(x):=\sigma_{t}F(x)+G(x)$ and $\widehat{\nabla\Phi}_{t}:=\sigma_{t}\widehat{\nabla F}_{t}+\widehat{\nabla G}_{t}$ to denote the gradient estimate at step $t$. We also denote $x_{\operatorname{opt}}\in X_{\operatorname{opt}}$ as an optimal solution to (1). All norms refer to the Euclidean norm.
For any sequence $\{x_{t}\}_{t\geq 0}$ generated by either Algorithm˜1 or Algorithm˜2, we observe that by construction, $x_{0},v_{t}\in X$ . Hence, under convexity of $X$, $x_{t+1}\in X$ as it is a convex combination of $x_{t}$ and $v_{t}$, and $X$ is assumed to be convex. By a simple induction argument, one can easily show that the sequence $\{x_{t}\}_{t\geq 0}\in X$.
A.1 Convex outer-level
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 or Algorithm˜2 with some given stepsizes $\{\alpha_{t}\}_{t\geq 0}$. If ˜1 holds, then for every $t\geq k$, we have
$\Phi_{t}(x_{t+1})-\Phi_{t}(x_{\operatorname{opt}})\!\leq\!(1-\alpha_{t})(\Phi_{t}(x_{t})-\Phi_{t}(x_{\operatorname{opt}}))+\frac{(L_{f}\sigma_{t}\!+\!L_{g})D^{2}\alpha_{t}^{2}}{2}+2\alpha_{t}D\left\|\nabla\Phi_{t}(x_{t})\!-\!\widehat{\nabla\Phi}_{t}\right\|,$
Proof of Lemma˜1.
@@
$x_{t+1}=x_{t}+\alpha_{t}(v_{t}-x_{t}),\quad\text{where}\quad v_{t}\in\operatorname*{arg\,min}_{v\in X}\left\{\left(\widehat{\nabla\Phi}_{t}\right)^{\top}v\right\}.$
@@
and set $w_{t}=x_{\operatorname{opt}}$. Thanks to the convexity of $\Phi_{t}$, one can easily show that $\nabla\Phi_{t}(x_{t})^{\top}(x_{\operatorname{opt}}-x_{t})\leq\Phi_{t}(x_{\operatorname{opt}})-\Phi_{t}(x_{t})$. Thus, we obtain
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 or Algorithm˜2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$. If ˜1 holds, then for each $t\geq k$, we have
@@
Applying Lemma˜1 with stepsize $\alpha_{t}=2/(t+2)$ for all $t\geq k$ gives the recursion $h_{t+1}\leq\frac{t}{t+2}h_{t}+\eta_{t}$ for any $t\geq k$. This can be re-expressed as
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 or Algorithm˜2 with stepsizes $\{\alpha_{t}\}_{t\geq 0}$ given by $\alpha_{t}=2/(t+2)$ for any $t\geq k$. If ˜1 holds, then for each $t\geq k$, we have
$\begin{aligned} G(x_{t})-G_{\operatorname{opt}}&\leq\frac{(k+1)k\left(\Phi_{k}(x_{k})-\Phi_{k}(x_{\operatorname{opt}})\right)}{t(t+1)}+\frac{2(\sigma_{0}L_{f}+L_{g})D^{2}(t-k)}{t(t+1)}\\[4.30554pt] &\quad+\sum_{i\in[t]\setminus[k]}\frac{4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|}{t(t+1)}\\[4.30554pt] &\quad+\left(F_{\operatorname{opt}}-\min_{x\in X}F(x)\right)\left(\sigma_{t}+\frac{\sum_{i\in[t]}(i+1)i(\sigma_{i-1}-\sigma_{i})}{t(t+1)}\right),\end{aligned}$ (10)
@@
Proof of Proposition˜1.
By definition of the function $\Phi_{t}$, the first term in Eq.˜9 satisfies
@@
Since $x_{t}\in X$, we have $\min_{x\in X}F(x)-F_{\operatorname{opt}}\leq F(x_{t})-F_{\operatorname{opt}}$ for all $t\geq 0$. Using this bound together with the identity in Eq.˜12, we can further lower bound the left-hand side of inequality (9), and arrive at the bound (10). This completes the proof of the first claim.
For the second claim, since $\{x_{t}\}_{t\geq 0}\in X$ is feasible in the lower-level problem, it follows directly that $G(x_{t})\geq G_{\operatorname{opt}}$. Combining this inequality with (12) and  (9), we obtain Eq.˜11, which concludes the proof. ∎
A.2 Nonconvex outer-level
When $F$ is possibly nonconvex, we start by establishing a bound, akin to Lemma 2.
@@ A.2 Nonconvex outer-level
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 or Algorithm˜2 with $\{\alpha_{t},\sigma_{t}\}_{t\geq 0}$ given by $\alpha_{t}=(t+1)^{-\omega}$ and $\sigma_{t}=\varsigma(t+1)^{-p}$ for $t\geq k$, with $0<p\leq\omega$. If ˜1 holds with the exception that $F$ may be nonconvex, then for every $t\geq k$, we have
Proof of Lemma˜3.
Using the definition of $\Phi_{t}=\sigma_{t}F+G$, for any $t\geq k$, we have
where the inequality follows directly from (8) and by minimizing over $w_{t}\in X$. We emphasize that (8) does not involve any convexity assumptions on $\Phi_{t}$ and relies solely on its smoothness. Furthermore, notice that for any $x\in X$,
$\leq\min_{v\in X_{\operatorname{opt}}}~\left\{\sigma\nabla F(x)^{\top}(v-x)+G(v)-G(x)\right\}$
$=-\sigma\mathcal{F}(x)-\mathcal{G}(x),$
$\begin{aligned} \alpha_{t}(\sigma_{t}\mathcal{F}(x_{t})+\mathcal{G}(x_{t}))&\leq\sigma_{t}(F(x_{t})-F(x_{t+1}))+\mathcal{G}(x_{t})-\mathcal{G}(x_{t+1})\\ &\quad+2D\alpha_{t}\|\widehat{\nabla\Phi}_{t}-\nabla\Phi_{t}(x_{t})\|+\frac{(\sigma_{t}L_{f}+L_{g})D^{2}\alpha_{t}^{2}}{2}.\end{aligned}$ (13)
Using the definition of $\alpha_{t}$ and the fact that $(t+1)^{\omega}\leq t^{\omega}-1$ for any $t\geq 0$, we arrive at
Unwinding this recursion back to $i=k$, we obtain
Since $p\leq\omega$, we have $(i+1)^{\omega}\sigma_{i}-i^{\omega}\sigma_{i-1}\geq 0$. Therefore, we deduce that
$\sum_{i=k}^{t}(i+1)^{\omega}\sigma_{i}(F(x_{i})-F(x_{i+1}))$
$\leq 2\varsigma\sup_{z\in X}|F(z)|(t+1)^{\omega-p}+\sigma_{k}F(x_{k}).$
This completes the proof. ∎
[…]
Lemma 4.
If ˜1 holds, the function $\mathcal{F}$ as defined in (3) is $L_{f}D+\max_{z\in X}\|\nabla F(z)\|$-Lipschitz continuous over $X$.
Proof of Lemma˜4.
Given $x,y\in X$, we observe that
$\mathcal{F}(y)$ $=\max_{v\in X_{\operatorname{opt}}}\left\{\nabla F(y)^{\top}(y-v)\right\}$
$=\max_{v\in X_{\operatorname{opt}}}\{(\nabla F(y)-\nabla F(x))^{\top}(y-v)+\nabla F(x)^{\top}(y-x)+\nabla F(x)^{\top}(x-v)\}$
$=\max_{v\in X_{\operatorname{opt}}}\{(\nabla F(y)-\nabla F(x))^{\top}(y-v)+\nabla F(x)^{\top}(y-x)+\nabla F(x)^{\top}(x-v)\}$
$\leq\max_{v\in X_{\operatorname{opt}}}\{\|\nabla F(y)-\nabla F(x)\|\|y-v\|\!+\nabla F(x)^{\top}(y-x)+\nabla F(x)^{\top}(x-v)\}$
$\leq\max_{v\in X_{\operatorname{opt}}}\{L_{f}\|x-y\|D+\nabla F(x)^{\top}(y-x)+\nabla F(x)^{\top}(x-v)\}$
$=L_{f}D\|x-y\|+\nabla F(x)^{\top}(y-x)+\mathcal{F}(x)$
$\leq\left(L_{f}D+\max_{z\in X}\|\nabla F(z)\|\right)\|y-x\|+\mathcal{F}(x),$
where the first and third inequality follows from Cauchy-Schwartz inequality, while the second inequality follows from the $L_{f}$-smoothness of $F$.
By interchanging the role of $y$ and $x$, we deduce that
$|\mathcal{F}(y)\!-\!\mathcal{F}(x)|\leq\left(L_{f}D+\max_{z\in X}\|\nabla F(z)\|\right)\|y-x\|.$
This concludes the proof. ∎
Appendix B Proofs of Section 2
B.1 Proof of Theorem˜1
We restate Theorem˜1 with exact upper bounds on sub-optimality gaps.
Theorem 1.
Let $\{z_{t}\}_{t\geq 1}$ be the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ given by $\sigma_{t}:=\varsigma(t+1)^{-p}$ for some chosen $\varsigma>0$, $p\in(0,1/2)$. Under ˜1, for any $t\geq 1$, with probability at least $1-\delta$, we have
$F(z_{t})-F_{\text{opt}}\leq\frac{2(L_{f}+L_{g}/\varsigma)D^{2}}{(t+1)^{1-p}}+\frac{4c\left(4(L_{f}+L_{g}/\varsigma)D+3(\sigma_{f}+\sigma_{g}/\varsigma)\right)}{(t+1)^{1/2-p}}\sqrt{\log\left({8dt^{2}}/{\delta}\right)},$
and
$G(z_{t})-G_{\text{opt}}$
$\leq\left(1+\frac{2p}{\min\{1,2(1-p)\}}\right)\left[(1+2p)\left(F_{\text{opt}}-\min_{x\in X}F(x)\right)\varsigma+2(\varsigma L_{f}+L_{g})D^{2}\right$.
$\quad\left.+4c\left(8\left(L_{f}\varsigma+L_{g}\right)D^{2}+3\left(\sigma_{f}\varsigma+\sigma_{g}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left({8d}/{\delta}\right)}\,\right](t+1)^{-p},$
where $c$ is some absolute constant. Moreover, with probability $1$, we have
The proof proceeds by applying Proposition˜1 with $k=0$. In this case, the main random component in Eq.˜10 and Eq.˜11 is the term $\sum_{i\in[t]}4iD\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$. We first derive a probabilistic upper bound for this term, then carefully set the parameters in Algorithm˜1 to establish the convergence rate for both the lower- and upper-level problems.
We begin by bounding the gradient estimator used in Algorithm˜1.
Lemma 5.
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with stepsize $\alpha_{t}=2/(t+2)$ for any $t\geq 0$. If ˜1 holds, then for any $t\geq 1$, given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$, it (jointly) holds that
@@ B.1 Proof of Theorem˜1
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$. If ˜1 holds and the parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing and positive, then given $\delta\in(0,1)$, with probability at least $1-\delta$, for some absolute constant $c$ and any $t\geq 1$, we have
$\sum_{i\in[t]}i\|\widehat{\nabla\Phi}_{i-1}-\nabla\Phi_{i-1}(x_{i-1})\|$
$\leq c\left(4(\sigma_{0}L_{f}+L_{g})D+3(\sigma_{0}\sigma_{f}+\sigma_{g})\right)(t+1)^{3/2}\sqrt{\log\left({8dt^{2}}/{\delta}\right)}.$
Proof of Proposition˜2.
Given $i\geq 1$ and $\delta\in(0,1)$, by Lemma˜5, with probability at least $1-\delta/i(i+1)$, we have
@@
To establish the desired convergence results for Algorithm˜1, we need to impose certain conditions on the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$, stated below.
@@
The regularization parameters $\{\sigma_{t}\}_{t\geq 0}$ are non-increasing, positive, and converge to 0.
@@
While it is not obvious a priori, Lemma˜17 in Section˜D.1 guarantees that these quantities remain finite for any parameter choice satisfying 1 to 3. In particular, 1 to 3 hold when $\sigma_{t}=\varsigma(t+1)^{-p}$ for any $p\in(0,1)$, which is what is used in Theorems˜1 and 3, and Lemma˜8 provides explicit bounds for this case. We now analyze the sequence $\{x_{t}\}_{t\geq 0}$ in Algorithm˜1 for the inner-level problem.
@@
Let $\{x_{t}\}_{t\geq 0}$ denote the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta}$ be defined as in (15a). If ˜1 and 1 to 3 hold, with probability at least $1-\delta$, it holds that
@@
Proof of Lemma˜6.
Combining (10) from Proposition˜1 with the probabilistic bound in Proposition˜2, and after some straightforward calculations, with probability at least $1-\delta$, we have
@@
We next analyze the convergence of the sequence $\{z_{t}\}_{t\geq 0}$ in Algorithm˜1 for both outer- and inner-level problems.
@@
Let $\{z_{t}\}_{t\geq 1}$ denote the iterates generated by Algorithm˜1 with $\alpha_{t}=2/(t+2)$ for any $t\geq 0$, and let $C_{\delta},V$ be defined as in Eq.˜15. If ˜1 and 1 to 3 hold, then with probability at least $1-\delta$, it (jointly) holds that
@@
Proof of Lemma˜7.
Recall from Algorithm˜1 that we have defined
@@
Using the inequality $S_{t}\geq(t+1)t\sigma_{t}$, which holds thanks to ˜1, and applying Proposition˜2, with probability at least $1-\delta$, we have
@@
Lemma˜7 establishes a convergence result in terms of the regularization parameters $\{\sigma_{t}\}_{t\geq 0}$. The next lemma specifies an update rule for these parameters and provides bounds on the quantities introduced in Eq.˜15.
@@
Consider the sequence $\sigma_{t}:=\varsigma(t+1)^{-p}$ for $t\geq 0$ and the quantities defined in Eq.˜15.
@@
    If $p\in(0,1)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies ˜1 and ˜2 with $L=p$. Furthermore,
@@
    If $p\in(0,1/2)$, then $\{\sigma_{t}\}_{t\geq 0}$ satisfies ˜3, and we have
    $\bar{C}_{\delta}\leq 4c\left(8\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left({8d}/{\delta}\right)}.$
Proof of Lemma˜8.
As for assertion (i), it is trivial to see that the sequence $\{\sigma_{t}\}_{t\geq 0}$ satisfies Conditions 1. To validate ˜2, observe that
@@
Dividing both sides by $t(t+1)\sigma_{t}^{2}$, we deduce that $V\leq 2p$. When 12p<01-2p<0, we have
@@
As for assertion (ii), it is trivial to see that ˜3 holds. We thus focus on bounding $\bar{C}_{\delta}$. Using the fact that $\sup_{t\geq 1}{\log(t)}/{t}={1}/{e},$ for any $t\geq 1$, we may conclude that
@@
$=8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}\!+\!3\left(\sigma_{f}\!+\!\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{1}{(t+1)^{1-2p}}\left(\frac{2}{1-2p}\log\left(t^{1-2p}\right)\!+\!\log\left({8d}/{\delta}\right)\right)}$
$\leq 8c\left(4\left(L_{f}+\frac{L_{g}}{\varsigma}\right)D^{2}+3\left(\sigma_{f}+\frac{\sigma_{g}}{\varsigma}\right)D\right)\sqrt{\frac{2}{(1-2p)e}+\log\left({8d}/{\delta}\right)}.$